- P04 INSTRUMENT HISTORY
- PRP INSTRUMENTS
- System shipped RSMAS to BNL for preparation for calibration.
- Note: condition from RSMAS was very bad. It had considerable soot/oil
contamination. The FRSR diffuser was very dirty. We cleaned it as much
as we could, but it still seems a little off color.
- This unit has an old 848 B&W pyranometer. It appears that water
and salt have entered the instrument. One can see salt crystals inside
the dome. On inspection, the o-ring for the desiccant plug seems to have
been broken and repaired by adding another o-ring on top of the old one.
This does not seal properly.
- Electronic re-calibration. Prpcal_04_0104 is almost identical
to PrpCal_04_med00 done a year ago.
- Deployed on Mt. Lemmon
- Returned to BNL. Heater circuit problems.
- At BNL, heater problem resolved.. Corrosion in the cable was found
right at pin 1 (the head thermistor.)
- Ship to RSMAS for Gulf of Mexico Cruise on R/V Justo Sierra.
Expected return 1 Aug.
- Changed shaft o-rings and the motor case cover. The cover now has a
"D" hole for the Conxall.
- Examined the erroneous head temperature readings. The first head temp.
reading taken by the A to D after the system goes into LOW MODE is bad while
the rest are normal. We've decided to leave this alone and continue to monitor
Re-calibrated the electronics in preparation of taking the unit to the
Akademic Ioffe in Ushoya. Installed on the roof.
Received P04 from RSMAS today for calibration and maintenance.
Sent head 462 to YES for calibration. Received a new PSP from RAMAS to replace
the 8-48 sensor that was there. Did a thermistor calibration on the PIR. Roof
tested for about 2 weeks. Shipped to Bob Evans on Thursday 08/29/02.
Received at BNL for maintenance.